OpenStudy (anonymous):

The number of real roots of (x+3)^4 + (x+5)^4 = 16 is: a) 0 b) 2 c) 4 d) none of these

OpenStudy (experimentx):

2

OpenStudy (experimentx):

-3 and -5

OpenStudy (anonymous):

Please say the detailed steps.

OpenStudy (experimentx):

here's an ugly method http://math.stackexchange.com/questions/785/is-there-a-general-formula-for-solving-4th-degree-equations simplify it to 4th degree equation ... it has at max 4 roots, in our case it's 2 real and 2 complex

OpenStudy (anonymous):

@experimentX can u explain the direct solution and step? I didn't get it.

OpenStudy (anonymous):

 the no of solution =cut points of graphs(y=(x+3)^4 and y=16-(x+5)^4)=2 (they are -3 and -5)